PX= $9500    PY= $10000  I = $15000     A = $170000  W = 160

This function is:

                       Qs = 89830 -40PS+20PX+15PY+2I +.001A +10W

Calculate the point price elasticity of demand at PS = $9000 (which should make QS = 101600). Does this elasticity (E) value indicate that Smooth Sailing boat demand is relatively responsive to Smooth Sailing boat price changes? Explain why or why not. The formula is E=(dQs/dPs)*(Ps/Qs)

Calculate the point weather elasticity of demand with W = 160. Use Qs corresponding to Ps = 9000.   Other variables and their values are as given at the top, before question #1. Does this elasticity indicate that the demand for Smooth Sailing's boats is relatively responsive to changes in the number of favorable weather days? Explain why or why not. The formula is E=(dQs/W)*(W/Qs)

Calculate the point income elasticity of demand, given that I = $170,000 and that PS = $8500 (thus QS should equal 431,600). Other variables are as given at the top before #1. Does this elasticity indicate that the demand for Smooth Sailing boats is relatively responsive to changes in income? Explain why or why not. The formula is E=(dQs/I)*(I/Qs)
Answer #1

Qs = 89830-40 ps + 20 P x + 15 Py+20 +0.001 A+low I= 15,000 At, PX 2 9500, A = 170000 PY=10,000 W = 160 Qs : 89830 - 40PS + 2

Inecome elastraly Now I 2 170,000 previon I = 15,000 SO, DI 170.000 - 15000 = 155000 1 so, Qs = 461600-40Ps + 2/155000) = 77